OpenStudy (anonymous):

In the function f(x) = sec 2x, what are the following: - Domain - Asymptotes

OpenStudy (solomonzelman):

Domain will have gaps. \(f(x)=\sec^2(x)\) \(f(x)=\displaystyle \frac{1}{ \cos^2(x)}\) so for every value where cos²(x)=0, you have a vertical asymptote.

OpenStudy (anonymous):

Wait, so \[\sec^2x \] is also \[\sec2x\]?

OpenStudy (anonymous):

Can I use the T-table to graph the function or nah?

OpenStudy (solomonzelman):

Also, you know that |cos x| < |sec x| And cos(x) (as well as sine), is between 1 and -1. This way, |sec x| is always ≥1 and ≤-1. So, it will look roughly like this: